OpenStudy (anonymous):

A cash register contains only five dollar and ten dollar bills. It contains twice as many five's as ten's and the total amount of money in the cash register is 580 dollars. How many ten's are in the cash register?

OpenStudy (dn_20):

So x = 5, y = 10, z= # of 5's have, and a= # of 10's have 2xz +ya =580 2 (5) z + 10a = 580 10z +10a = 580 10(z+a) = 580 z+a =58 We have twice as many fives than tens, so z = 2a Substitution Property of Equality 2a +a = 58 3a = 58 a = 19.333 So z= 38.667 So we have 19 bill of 10's and 38 bill of 5's, but the last one? 1/3 for 10's and 2/3 for 5's, so 19 1/3 bill of 10's and 38 2/3 bill of 5's

OpenStudy (dn_20):

Sorry, I not totally sure if I did it right. But I hope that I'm right since I checked many time before I sent this.

OpenStudy (anonymous):

what if the money was 680 would it be the same formula

OpenStudy (dn_20):

Yeah

OpenStudy (dn_20):

The answers is crazy, since the bills cannot be in fraction

OpenStudy (anonymous):

So if its 680, how many 10s would there be?

OpenStudy (dn_20):

The number of 10s will be 22.667, while the 5s are 45.333